OpenStudy (anonymous):

What is Z? 2+ 1/5z=7

OpenStudy (anonymous):

\[2+\frac{ 1 }{5z }=7\] multiply both sides by 5z to cancel remove the fraction... \[\left( 2+\frac{ 1 }{5z }=7 \right)\times5z\] \[10z+1=35z\] did you get it up to here? then combine all the terms that are similar( or has z on one side of the equation)

OpenStudy (bookworm14):

7-2 = 5 so now you have to solve for 1/5 ..... 1/5 * z = 5 ... z = 5 2+1/5*5 = 7